# Understanding How Psychological Distance Influences User Preferences in Conversational Versus Web Search

### v1

@@ 3.1. Task Design
In the experiments, participants were asked to take an information seeking task with fictional travel scenarios. In the tasks, participants were asked to imagine planning a one-week vacation to another country, focusing on food, natural attractions, and historical buildings.
We varied the psychological distance between the participant and the task as the main independent variable (two levels: near or far) in a between-subjects design, and repeated this across four dimensions of psychological distance (spatial, temporal, social, or hypothetical). These designs allowed us to explore whether there are differences in user perceptions and preferences towards IR technologies based on psychological distance, and whether these differences are consistent across the different dimensions or influenced by other factors.
@@
Neutral Nearby country Next month The participant themselves (Not mentioned)
We varied a particular aspect of the tasks  [^1] Details on the specific scenarios are fully outlined in Appendix 7.1. according to the psychological dimension being manipulated in each experimental group, as seen in Table 1. We introduced a ”neutral” setting in our study, which does not constitute an independent experimental condition. Instead, it ensures that dimensions of psychological distance other than the one currently under focus remain neutral. This prevents these dimensions from influencing participants’ perception of the psychological distance being studied. For example, in ”neutral” distances, the choice of a nearby country of the residence of our participants as a destination reflects the reality that most international travel typically involves neighboring countries (McKercher and Mak, 2019; McKercher et al., 2008). Similarly, in describing the traveler in dimensions other than social distance, we choose the participant themselves because conducting information searches and planning trips for oneself is more common in everyday life.
Each participant was required to use both a WebSearch and a ConvSearch to perform the same tasks for different destinations, so that the first search using WebSearch did not influence the subsequent search with ConvSearch. In the same experimental group, the difference in destination did not affect psychological distance: Nearby countries in the set of near or neutral spatial distances were represented by Mexico (Information Seeking Task 1) and Cuba (Information Seeking Task 2), both geographically close to our participants in the United States, while faraway countries in the set of far spatial distances were represented by Egypt (Information Seeking Task 1) and Sudan (Information Seeking Task 2). These countries were chosen to minimize the influence of prior knowledge by excluding mainstream American travel destinations and focusing on non-Western countries.
@@ 3.2. Procedure
This figure presents a detailed flowchart of the study procedure, organized into six main phases across a horizontal layout. Each phase is represented as a rectangular box containing lists of activities or survey components, with arrows indicating the sequence from left to right. Step 1 is Pre Task Survey: This initial phase assesses participants’ baseline attitudes AI. Step 2 is Information Seeking Task 1: This phase starts with reading a scenario, followed by using the WebSearch to gather information. Participants’ search behaviors are recorded, and they are required to complete a travel plan. Step 3 is Post Task 1 Survey: This survey evaluates participants’ perceptions after the first task, focusing on Perceived Ease of Use, Usefulness, Enjoyment, and Information Credibility. Step 4 is Information Seeking Task 2: Similar to the first task, participants read a new scenario and use a Conversational Search. This phase also involves recording the conversation and completing a travel plan. Step 5 is Post Task 2 Survey: This survey mirrors the first post-task survey but evaluates the conversational search experience, assessing the same four perception metrics. Preference Survey: The final phase involves a survey where participants choose their preferred information retrieval system between WebSearch and the conversational search. They also provide reasons for their choice.
@@
The full study design (Figure 2) includes six phases. In Step 1, participants completed a pre-task survey to measure their baseline attitudes to artificial intelligence (AI). In Step 2, the scenario for the first information seeking task was introduced. Participants then used WebSearch to seek information. We placed WebSearch first as it is a widely recognized IR system, and serves as a baseline for the experiment. To monitor the search behaviors of participants, we created a custom web search interface that integrated Google’s search API  [^2] https://developers.google.com/custom-search/v1/overview. After gathering the necessary information for the task, participants were required to write a travel plan consisting of at least 100 words to ensure the task completion quality. Step 3 was the first post-task survey, where participants were asked to rate their perceptions of ease of use, usefulness, enjoyment, and information credibility when using WebSearch. Steps 4 and 5 mirrored the previous two steps. Participants used ConvSearch to perform the second information seeking task, which had a different destination that had the same spatial distance as the first one. We built the ConvSearch system on the UChat [^3] https://uchat.au/ platform using OpenAI’s GPT-4 API  [^4] https://platform.openai.com/docs/api-reference/completions/object, similar to ChatGPT. Participants used our tool to seek information and then reported their perceptions of this technology in the second post-task survey. Finally, in Step 6, participants indicated their preference between the two IR systems they used and elaborated on reasons for their choice.
@@ 3.3. Participants
Participants were recruited on the CloudResearch Connect platform [^5] https://www.cloudresearch.com/products/connect-for-researchers/, with residency restricted to the United States to ensure uniformity in their perception of spatial distances to designated travel destinations. After excluding individuals who failed attention checks, the study consisted of 128 participants according to the power analysis result (16 per group; 62 females, 64 males, 2 undisclosed). The average age of the participants was 37.6 (SD = 10.5), and 63% of the participants had at least a bachelor’s degree. The study was expected to last 30 minutes, with participants compensated at CloudResearch’s recommended rate of \$10 per hour. This research was reviewed and approved by the Anonymous University Departmental Ethics Review Committee.
@@ 3.4.1. User perception, behavior and preferences
3.4.2. Thematic Analysis About User Preferences Reasons
We conducted a thematic analysis (Braun and Clarke, 2006) of the reasons participants (n = 128) reported their tool preferences to provide user-centered design recommendations for IR tools. Initially, two researchers independently performed semantic (what the participants said) and latent (researchers’ interpretations of the data) analyses on the data to generate preliminary codes (without predefined categories). The research team refined these initial codes through multiple discussions and developed themes using a bottom-up inductive approach. Subsequently, two researchers coded the reasons participants reported, potentially assigning multiple themes to each reason.
@@ 3.4.3. Control Variables
The Pearson correlation coefficients between the control variable and all dependent variables did not reach statistical significance (all p-values ¿ 0.05)  [^6] Specific data are presented in Appendix 7.2..
@@ 4. Results
4.1. Greater Psychological Distance Increases Relative User Perceptions of Conversational Search
Our results generally show that as psychological distance increases from near to far, relative ratings of conversational search (i.e. ConvSearch rating - WebSearch rating) across multiple metrics increased. At near psychological distances, users were generally indifferent between the two systems; at far distances, users instead mostly preferred ConvSearch.
4.1.1. Partial Increase in Relative Perceived Ease of Use (H1 Partially Supported)
Figure 3. Comparison of perceived ease of use between WebSearch and ConvSearch In Far Psychological Distances. Panels (a)-(d) show results at a near psychological distance, and panels (e)-(h) at a far psychological distance, highlighting significant differences in the latter. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@ 4.1.1. Partial Increase in Relative Perceived Ease of Use (H1 Partially Supported)
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived ease of use for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived ease of use at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.490), marked as non-significant (ns); (b) Temporal Distance (p = 0.270), marked as non-significant (ns); (c) Social Distance (p = 0.153), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.560), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived ease of use at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. Panel (e) Spatial Distance shows a p-value of (p = 0.095), marked as non-significant (ns). The p-values indicate significant differences in panels (f) Temporal Distance (p = 0.025), marked with one asterisk (*); (g) Social Distance (p = 0.003), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.004), marked with two asterisks (**).
@@
When psychological distance was near, participants reported similar scores of perceived ease of use for WebSearch (Spatial Distance: $M=6.19$, $SD=0.96$; Temporal Distance: $M=5.91$, $SD=0.79$; Social Distance: $M=5.63$, $SD=1.32$; Hypothetical Distance: $M=5.81$, $SD=1.48$) and ConvSearch (Spatial Distance: $M=5.91$, $SD=1.35$; Temporal Distance: $M=5.55$, $SD=1.22$; Social Distance: $M=6.01$, $SD=0.97$; Hypothetical Distance: $M=6.03$, $SD=1.09$). As depicted in Fig. 3 (a)-(d), paired samples T-Tests showed non-significant differences with low effect sizes (Spatial Distance: $t(15)=0.71$, $p=0.49$, Cohen’s $d=0.18$; Temporal Distance: $t(15)=1.15$, $p=0.27$, Cohen’s $d=0.29$; Social Distance: $t(15)=-1.51$, $p=0.15$, Cohen’s $d=-0.38$; Hypothetical Distance: $t(15)=-0.60$, $p=0.56$, Cohen’s $d=-0.15$).
@@
In contrast, there were significant differences in perceived ease of use at a far psychological distance in three out of the four dimensions (Fig. 3 (f)-(h)). For temporal distance, ConvSearch demonstrated a higher ease of use ($M=6.18$, $SD=0.73$) compared to WebSearch ($M=5.65$, $SD=1.09$), with significant results, $t(15)=-2.14$, $p=\textbf{0.025}$, Cohen’s $d=-0.53$. Similarly, for social distance, ConvSearch ($M=6.22$, $SD=0.74$) outperformed WebSearch ($M=5.26$, $SD=1.10$), $t(15)=-3.14$, $p=\textbf{0.003}$, Cohen’s $d=-0.78$, and for hypothetical distance, ConvSearch ($M=6.30$, $SD=0.76$) was rated higher than WebSearch ($M=5.40$, $SD=1.23$), $t(15)=-3.06$, $p=\textbf{0.004}$, Cohen’s $d=-0.77$. However, with spatial distance (Fig. 3 (e)), although ConvSearch scored higher ($M=6.28$, $SD=0.48$) compared to WebSearch ($M=6.02$, $SD=0.73$), the difference was not statistically significant, $t(15)=-1.37$, $p=0.095$, Cohen’s $d=-0.34$. These findings mostly support H1, indicating that as psychological distance increases, ConvSearch tends to be perceived as easier to use compared to WebSearch, particularly in the temporal, social, and hypothetical dimensions.
@@
4.1.2. Increase in Relative Perceived Information Credibility (H2 Supported)
@@
Figure 4. Impact of psychological distance on perceived information credibility for WebSearch and ConvSearch. Panels (a)-(d) illustrate results at a near psychological distance, with significant differences only in spatial distance. Panels (e)-(h) demonstrate increased credibility for ConvSearch at a far psychological distance across all dimensions except temporal distance. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived information credibility for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived information credibility at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.011), marked with one asterisk (*); (b) Temporal Distance (p = 0.131), marked as non-significant (ns); (c) Social Distance (p = 0.612), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.550), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived information credibility at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.001), marked with three asterisks (***); panels (f) Temporal Distance (p = 0.012), marked with one asterisk (*); (g) Social Distance (p ¡ 0.001), marked with three asterisks (***); (h) Hypothetical Distance (p ¡ 0.001), marked with three asterisks (***).
At a near psychological distance, as depicted in Fig. 4 (b)-(d), participants reported similar perceptions of information credibility in three out of four dimensions between WebSearch (Temporal: $M=76.88$, $SD=15.86$; Social: $M=80.69$, $SD=18.37$; Hypothetical: $M=75.40$, $SD=25.22$) and ConvSearch (Temporal: $M=70.81$, $SD=21.86$; Social: $M=78.88$, $SD=22.58$; Hypothetical: $M=74.40$, $SD=27.80$). These differences were not statistically significant (Temporal: $t(15)=1.60$, $p=0.13$, Cohen’s $d=0.40$; Social: $t(15)=0.52$, $p=0.61$, Cohen’s $d=0.13$; Hypothetical: $t(15)=0.61$, $p=0.55$, Cohen’s $d=0.15$). However, for spatial distance (Fig. 4 (a)), ConvSearch ($M=75.46$, $SD=15.63$) demonstrated a significant credibility advantage over WebSearch ($M=59.28$, $SD=16.38$), $t(15)=-2.90$, $p=0.011$, Cohen’s $d=-0.73$.
@@
In contrast, when psychological distance was far, the perceived credibility of information from ConvSearch was significantly higher compared to WebSearch in all four dimensions (Fig. 4 (e)-(h)). For spatial distance, ConvSearch was rated as more credible ($M=76.38$, $SD=13.28$) than WebSearch ($M=56.88$, $SD=20.03$), yielding a significant difference, $t(15)=-4.92$, $p<\textbf{0.001}$, Cohen’s $d=-1.23$. This was also true for temporal distance (($M=77.94$, $SD=13.36$) vs. ($M=71.21$, $SD=18.73$), $t(15)=-2.50$, $p=\textbf{0.012}$, Cohen’s $d=-0.62$), social distance ($M=85.19$, $SD=13.11$) vs. ($M=70.31$, $SD=20.38$), $t(15)=-3.98$, $p<\textbf{0.001}$, Cohen’s $d=-1.00$), and hypothetical distance ($M=87.54$, $SD=10.11$) vs. ($M=74.23$, $SD=11.37$), $t(15)=-4.12$, $p<\textbf{0.001}$, Cohen’s $d=-1.03$. These findings support H2. They suggest that as psychological distance increases, so does the perceived reliability of information from ConvSearch compared to WebSearch.
4.1.3. Increase in Relative Perceived Usefulness (H3 Supported)

### v2

@@ 4.1. Task Design
Participants were asked to complete an information seeking task with fictional travel scenarios. The tasks required participants to imagine planning a one-week vacation to another country, focusing on finding food, natural attractions, and historical buildings destinations.
Each participant was randomly presented with one of eight varying tasks. Each task was either near or far in psychological distance to the participant, and psychological distance was manipulated in one of four dimensions (spatial, temporal, social, or hypothetical). For instance, a participant in the (near, spatial) condition was asked to plan a trip to a nearby country, while a participant in the (far, temporal) condition planned a trip in the distant future. This allowed us to explore whether there are differences in user perceptions and preferences towards IR technologies based on psychological distance, and whether these differences are consistent across the different dimensions or are influenced by other factors. Table 1 describes the ways in which all the different dimensions and distances were manipulated. [^1] For each task, dimensions that were not varied were given a ”neutral” setting. For example, in the (near, spatial) condition, the social, temporal, and hypothetical dimensions were kept neutral. These were to ensure that dimensions of psychological distance other than the one currently under focus remained neutral. In determining the neutral conditions, we aimed to prevent these other dimensions from influencing participants’ perception of the psychological distance being studied. For example, the choice of a destination country near to the residence of our participants reflects the reality that most international travel typically involves neighboring countries (McKercher and Mak, 2019; McKercher et al., 2008). Similarly, in describing the traveler in dimensions other than social distance, we chose the participant themselves because conducting information searches and planning trips for oneself is more common in everyday life.  [^2] Details on the specific scenarios are given in full in the Appendix (A.1), uploaded with Supplementary Material..
Each participant performed the same tasks but with different destinations for WebSearch and ConvSearch, so that the first search using WebSearch did not influence the subsequent search with ConvSearch. In each experimental group, the difference in destination did not affect psychological distance: Nearby countries in the set of near spatial distances were designated as Mexico (WebSearch task) and Cuba (ConvSearch task), both geographically close to our participants in the United States, while countries in the far spatial distance settings were represented by Egypt (WebSearch task) and Sudan (ConvSearch task). These countries were chosen to minimize the influence of prior knowledge by excluding mainstream American travel destinations and focusing on non-Western countries.
@@ 4.2. Procedure
This figure presents a detailed flowchart of the study procedure, organized into six main phases across a horizontal layout. Each phase is represented as a rectangular box containing lists of activities or survey components, with arrows indicating the sequence from left to right. Step 1 is Pre Task Survey: This initial phase assesses participants’ baseline attitudes towards AI. Step 2 is Information Seeking Task 1: This phase starts with reading a scenario, followed by using the WebSearch to gather information. Participants’ search behaviors are recorded, and they are required to complete a travel plan. Step 3 is Post Task 1 Survey: This survey evaluates participants’ perceptions after the first task, focusing on Perceived Ease of Use, Usefulness, Enjoyment, and Information Credibility. Step 4 is Information Seeking Task 2: Similar to the first task, participants read a new scenario and use a Conversational Search. This phase also involves recording the conversation and completing a travel plan. Step 5 is Post Task 2 Survey: This survey mirrors the first post-task survey but evaluates the conversational search experience, assessing the same four perception metrics. Preference Survey: The final phase involves a survey where participants choose their preferred information retrieval system between WebSearch and the conversational search. They also provide reasons for their choice.
@@
The full study design (Figure 2) includes six phases. In Step 1, participants completed a pre-task survey to measure their baseline attitudes to artificial intelligence (AI). In Step 2, the scenario for the first information seeking task was introduced. Participants then used WebSearch to seek information. We placed WebSearch first as it is a widely recognized IR system, and serves as a baseline for the experiment. To monitor the search behaviors of participants, we created a custom web search interface that integrated Google’s search API  [^3] https://developers.google.com/custom-search/v1/overview. After gathering the necessary information for the task, participants were required to write a travel plan consisting of at least 100 words to ensure the task completion quality. Step 3 was the first post-task survey, where participants were asked to rate their perceptions of ease of use, usefulness, enjoyment, and information credibility when using WebSearch. Steps 4 and 5 mirrored the previous two steps. Participants used ConvSearch to perform the second information seeking task, which had a different destination that had the same spatial distance as the first one. We built the ConvSearch system on the UChat [^4] https://uchat.au/ platform using OpenAI’s GPT-4 API  [^5] https://platform.openai.com/docs/api-reference/completions/object, similar to ChatGPT. Participants used our tool to seek information and then reported their perceptions of this technology in the second post-task survey. To minimize cognitive load and ensure consistency in participants’ experiences, we intentionally placed WebSearch first. It serves as a baseline for the experiment, as it is a system that participants are more familiar with in their daily lives. Previous studies have shown that using a system with which users are already familiar can reduce the cognitive load associated with the task (Blackler et al., 2010; Aarts et al., 1998; Norman and Bobrow, 1975), ensuring that users complete Steps 4 and 5 in a more natural state. Furthermore, as our study focuses on the relative comparisons between WebSearch and ConvSearch under different psychological distances, any potential bias introduced by the fixed order is consistent across all conditions and unlikely to confound the key findings. Finally, presenting the less familiar system first could have introduced additional variability, as participants might require more time to adapt to the new interface, which could inadvertently affect their perceptions (Limayenm et al., 2003). Finally, in Step 6, participants indicated their preference between the two IR systems they used and elaborated on reasons for their choice.
@@ 4.3. Participants
Participants were recruited on the CloudResearch Connect platform [^6] https://www.cloudresearch.com/products/connect-for-researchers/, with residency restricted to the United States to ensure uniformity in their perception of spatial distances to designated travel destinations. After excluding individuals who failed attention checks, the study consisted of 128 participants according to the power analysis result (16 per group; 62 females, 64 males, 2 undisclosed). The average age of the participants was 37.6 (SD = 10.5), and 63% of the participants had at least a bachelor’s degree. The study was expected to last 30 minutes, with participants compensated at CloudResearch’s recommended rate of \$10 per hour. This research was reviewed and approved by the National University of Singapore Departmental Ethics Review Committee.
@@ 4.4.2. Control Variables
While attitudes toward AI are not the primary variables we observe, they could potentially influence participants’ perceptions of ConvSearch. Biases or preconceived notions about AI might affect how individuals evaluate and utilize IR systems. For example, participants with a positive attitude toward AI may be more inclined to perceive these tools as useful and trustworthy. Therefore, we consider participants’ attitudes toward AI as a control variable to analyze its linear relationship with our dependent variables. This approach ensures that our analysis accurately reflects the impact of the tools themselves, without being confounded by individual biases toward artificial intelligence.
4.5. Data Analysis
To mitigate the influence of control variables, we first analyzed the relationship between participants’ overall willingness towards AI and the primary observed variables using the Pearson correlation coefficient (Pearson, 1895)  [^7] Specific data are presented in the Appendix (A.2), uploaded with Supplementary Material..
To test our hypotheses, we conducted paired sample t-tests (Student, 1908) on the dependent variables (Ease of Use, Information Credibility, Usefulness, and Enjoyment). This method allows us to compare participants’ perceptions of the same variables after using both WebSearch and ConvSearch in the same experiment. Initially, we assessed the normality of the observed variables using the Shapiro-Wilk test (Shapiro and Wilk, 1965)  [^8] Detailed Shapiro-Wilk test results for each dependent variable and the type of paired samples t-test used are available in the Appendix (A.5), uploaded with Supplementary Material.. For datasets that followed a normal distribution, we applied the student’s paired sample t-test to evaluate differences between conditions. If the normality assumption was violated, we used the Wilcoxon signed-rank test (Wilcoxon, 1992) as a non-parametric alternative. To control the False Discovery Rate (FDR) across multiple tests, we applied the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995) as a correction for multiple comparisons to the 32 paired sample t-tests conducted across the four dependent variables and psychological distance dimensions. Adjusted p-values are reported unless otherwise noted. For the binary data on user preference, we employed the chi-square test (Pearson, 1900) to examine changes in preference for ConvSearch under different psychological distances.
To further address our research questions, we conducted a thematic analysis (Braun and Clarke, 2006) of the reasons participants (n = 128) reported for their tool preferences, aiming to provide user-centered design recommendations for IR tools. Initially, two researchers independently performed semantic (what the participants said) and latent (researchers’ interpretations) analyses to generate preliminary codes without predefined categories. The research team refined these codes through multiple discussions and developed themes using a bottom-up inductive approach. Subsequently, two researchers coded the reasons participants provided, potentially assigning multiple themes to each reason.
@@ 5. Results
5.1. Psychological Distance Increases Differences in User Perceptions
5.1.1. The Difference in Perceived Ease of Use Increased (H1 Partially Supported)
Figure 3. Comparison of perceived ease of use between WebSearch and ConvSearch in Far Psychological Distances. Panels (a)-(d) show results at a near psychological distance, and panels (e)-(h) at a far psychological distance, highlighting significant differences in the latter. Significance is marked as p ¿ 0.05 (ns), p ¡= 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@ 5.1.1. The Difference in Perceived Ease of Use Increased (H1 Partially Supported)
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived ease of use for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks from the paired samples statistical tests. In the first row, panels (a) through (d) depict the perceived ease of use at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.689), marked as non-significant (ns); (b) Temporal Distance (p = 0.360), marked as non-significant (ns); (c) Social Distance (p = 0.233), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.964), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived ease of use at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. Panel (e) Spatial Distance shows a p-value of (p = 0.163), marked as non-significant (ns). The p-values indicate significant differences in panels (f) Temporal Distance (p = 0.050), marked with one asterisk (*); (g) Social Distance (p = 0.004), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.004), marked with two asterisks (**).
@@
When psychological distance was near, participants reported similar scores of perceived ease of use for WebSearch (Spatial Distance: $M=6.19$, $SD=0.96$; Temporal Distance: $M=5.91$, $SD=0.79$; Social Distance: $M=5.63$, $SD=1.32$; Hypothetical Distance: $M=5.81$, $SD=1.48$) and ConvSearch (Spatial Distance: $M=5.91$, $SD=1.35$; Temporal Distance: $M=5.55$, $SD=1.22$; Social Distance: $M=6.01$, $SD=0.97$; Hypothetical Distance: $M=6.03$, $SD=1.09$). As depicted in Fig. 3 (a)-(d), paired samples statistical tests showed non-significant differences with low effect sizes (Spatial Distance: $W=32.50$, $p=0.689$, $r=0.18$; Temporal Distance: $t(15)=1.15$, $p=0.360$, Cohen’s $d=0.29$; Social Distance: $t(15)=-1.51$, $p=0.233$, Cohen’s $d=-0.38$; Hypothetical Distance: $W=32.00$, $p=0.964$, $r=-0.03$). The positive effect sizes for spatial and temporal distance suggest that when information seeking tasks are framed with near spatial or temporal distances, participants perceived WebSearch as slightly easier to use compared to ConvSearch. Conversely, the negative effect sizes for Social Distance and Hypothetical Distance indicate that when these tasks are framed with near social or hypothetical distances, participants perceived ConvSearch as slightly easier to use compared to WebSearch. However, the differences in perceived ease of use between the two tools under near psychological distances were not statistically significant and are negligible in practical terms.
@@
In the context, as shown in Fig. 3 (f)-(h), significant differences emerged in the perceived ease of use at a far psychological distance in the temporal, social, and hypothetical dimensions with a moderate or large effect size. Specifically, for temporal distance, ConvSearch demonstrated a significantly higher ease of use ($M=6.18$, $SD=0.73$) compared to WebSearch ($M=5.65$, $SD=1.09$), with a moderate effect size ($t(15)=-2.14$, $p=0.050$, Cohen’s $d=-0.53$). Similarly, in social distance, ConvSearch ($M=6.22$, $SD=0.74$) outperformed WebSearch ($M=5.26$, $SD=1.10$), with a large effect size ($W=2.00$, $p=0.004$, $r=-0.96$). Lastly, in the hypothetical distance dimension, ConvSearch ($M=6.30$, $SD=0.76$) was rated higher than WebSearch ($M=5.40$, $SD=1.23$), also with a large effect size ($W=2.00$, $p=0.004$, $r=-0.96$). For spatial distance (Fig. 3 (e)), although ConvSearch ($M=6.28$, $SD=0.48$) scored higher than WebSearch ($M=6.02$, $SD=0.73$), the difference was not statistically significant, $W=8.00$, $p=0.163$, $r=-0.56$. However, the moderate effect size suggests a meaningful trend favoring ConvSearch under far spatial distances. The positive effect sizes observed for far temporal, social, and hypothetical distances indicate that as the psychological distance of tasks increases, participants increasingly perceive ConvSearch as easier to use compared to WebSearch. These findings partially support H1, indicating that as psychological distance increases, ConvSearch tends to be perceived as easier to use compared to WebSearch, particularly in temporal, social, and hypothetical dimensions.
@@
5.1.2. Psychological Distance Increases Differences in Perceived Information Credibility (H2 Supported)
@@
Figure 4. Impact of psychological distance on perceived information credibility for WebSearch and ConvSearch. Panels (a)-(d) illustrate results at a near psychological distance, with significant differences only in spatial distance. Panels (e)-(h) demonstrate increased credibility for ConvSearch at a far psychological distance across all dimensions except temporal distance. Significance is marked as p ¿ 0.05 (ns), p ¡= 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived information credibility for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks from the paired samples statistical tests. In the first row, panels (a) through (d) depict the perceived information credibility at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.210), marked as non-significant (ns); (b) Temporal Distance (p = 0.210), marked as non-significant (ns); (c) Social Distance (p = 0.386), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.607), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived information credibility at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.044), marked with one asterisk (*); panels (f) Temporal Distance (p = 0.033), marked with one asterisk (*); (g) Social Distance (p ¡ 0.001), marked with three asterisks (***); (h) Hypothetical Distance (p = 0.004), marked with two asterisks (**).
At a near psychological distance, as depicted in Fig. 4 (a)-(d), participants reported similar perceptions of information credibility related to travel between WebSearch (Spatial Distance: $M=83.02$, $SD=14.47$; Temporal Distance: $M=76.88$, $SD=15.86$; Social Distance: $M=80.69$, $SD=18.37$; Hypothetical Distance: $M=75.40$, $SD=25.22$) and ConvSearch (Spatial Distance: $M=75.46$, $SD=15.63$; Temporal Distance: $M=70.81$, $SD=21.86$; Social Distance: $M=78.88$, $SD=22.58$; Hypothetical Distance: $M=74.40$, $SD=27.80$). Paired samples statistical tests showed non-significant differences with small effect sizes across dimensions (Spatial Distance: $t(15)=1.62$, $p=0.210$, Cohen’s $d=0.41$; Temporal Distance: $t(15)=1.60$, $p=0.210$, Cohen’s $d=0.40$; Social Distance: $W=42.00$, $p=0.386$, $r=-0.30$; Hypothetical Distance: $t(15)=0.61$, $p=0.607$, Cohen’s $d=0.15$). The small effect sizes observed at near distances indicate that psychological proximity of the task does not strongly influence users’ perceived credibility between the two search modalities. Notably, the slightly positive effect sizes for spatial, temporal, and hypothetical distances suggest a marginal preference for WebSearch in terms of credibility at near distances. Conversely, the negative effect size for social distance implies a slight preference for ConvSearch in this context.
@@
In the context of greater psychological distance, the perceived credibility of information from ConvSearch was significantly higher compared to WebSearch across all measured dimensions, as shown in Fig. 4 (e)-(h). For spatial distance, ConvSearch was rated as more credible ($M=76.38$, $SD=13.28$) than WebSearch ($M=70.67$, $SD=19.50$), yielding a significant difference with a moderate effect size, $t(15)=-2.23$, $p=0.044$, Cohen’s $d=-0.56$. In the temporal distance dimension, ConvSearch also demonstrated significantly higher credibility ($M=77.94$, $SD=13.36$) compared to WebSearch ($M=71.21$, $SD=18.73$), with a moderate effect size, $t(15)=-2.50$, $p=0.033$, Cohen’s $d=-0.62$. Similarly, in the social distance dimension, participants rated the credibility of ConvSearch ($M=85.19$, $SD=13.11$) significantly higher than that of WebSearch ($M=70.31$, $SD=20.38$), with a large effect size ($W=0.00$, $p<0.001$, $r=-1.00$). Lastly, for hypothetical distance, ConvSearch ($M=87.54$, $SD=10.11$) was perceived as more credible than WebSearch ($M=74.23$, $SD=11.37$), with a large effect size ($t(15)=-4.12$, $p=0.004$, Cohen’s $d=-1.03$). The moderate and large effect sizes observed at far psychological distances suggest that as the psychological distance of a task increases, ConvSearch becomes substantially more credible in the eyes of users compared to WebSearch. The negative direction of the effect sizes consistently indicates that participants perceived ConvSearch as more credible than WebSearch across all far-distance dimensions. These findings support H2. They suggest that as psychological distance increases, so does the perceived reliability of information from ConvSearch compared to WebSearch, even when the IR technologies remain unchanged. This indicates that psychological distance can influence user perceptions of information credibility in different contexts.
5.1.3. Psychological Distance Increases Differences in Perceived Usefulness (H3 Supported)